OpenStudy (anonymous):

find the slope that passes through the points (6/7,1 1 /5) and (4/7, 2/5)

OpenStudy (anonymous):

The slope is the rise divided by the run. The run is 6/7-4/7= 2/7 The rise is 6/5-2/5= 4/5 So divide 4/5 by 2/7= 4/5 * 7/2= 28/10 or 2 4/5